# On De Giorgi’s lemma for variational interpolants

### v2

@@ 2.1 Metric energy identity and De Giorgi’s estimate: statements and examples
Under Hypothesis 2.5, any measurable selection $(0,\sigma_{*})\ni\sigma\to\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ fulfills
${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma\,\psi\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}+\int_{0}^{\sigma}\psi^{*}\big{(}\psi^{\prime}\big{(}\frac{1}{\rho}\mathsf{d}^{\pm}_{\rho}(u^{o})\big{)}\big{)}\;\!\mathrm{d}\rho={\mathcal{E}}(u^{o}).$ (2.11)
There is a straightforward way to relate the metric energy identity to De Giorgi’s estimate, and that is throughout the following result in which the slope at $\widetilde{u}_{\sigma}$ is estimated in terms of the slope of the distance function ${\mathcal{D}}(u^{o},\cdot)$. In fact, (2.12) below extends to the setup of a $\mathrm{gMGS}$ $(M,{\mathcal{E}},{\mathcal{D}},\psi)$, the slope estimate proved in [AGS05, Lem. 3.1.3] in the quadratic case $\psi(r)=\frac{1}{2}r^{2}$. For completeness, we also record that a version of (2.12) was proved in [RMS08, Lemma 4.4] for non-differentiable dissipation potentials $\psi$. Note that this estimate is the metric counterpart of the Euler-Lagrange equation in the Banach-space setting, and hence it is less precise for a general $\mathrm{gMGS}$.
@@
$\big{|}\partial{\mathcal{E}}\big{|}(\widetilde{u}_{\sigma})\leq\psi^{\prime}\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}\;\big{|}\partial({-}{\mathcal{D}})(u^{o},\cdot)\big{|}(\widetilde{u}_{\sigma})\leq\psi^{\prime}\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}.$ (2.12)
@@
${\mathcal{E}}(\widetilde{u}_{\sigma})-{\mathcal{E}}(v)$ $=\Phi_{\sigma}(u^{o},\widetilde{u}_{\sigma})-\Phi_{\sigma}(u^{o},v)-\sigma\psi\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}+\psi\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},v)\big{)}$
$\stackrel{{\scriptstyle(1)}}{{\leq}}\sigma\Big{(}\psi\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},v)\big{)}-\psi\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}\Big{)}$
$\stackrel{{\scriptstyle(2)}}{{\leq}}\psi^{\prime}\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}\>\Big{(}{-}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})+{\mathcal{D}}(u^{o},v)\Big{)}.$
where (1) is due to $\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$, whereas (2) follows by the convexity of $\psi$. Taking the positive part on both sides (using $\psi^{\prime}\geq 0$), dividing by ${\mathcal{D}}(\widetilde{u}_{\sigma},v)$, and taking the limsup for $v\to\widetilde{u}_{\sigma}$ gives the first estimate in (2.12).
By the triangle inequality for the distance ${\mathcal{D}}$, for any fixed $u^{0}\in M$ the functions ${\mathcal{D}}(u^{0},\cdot)$ and ${-}{\mathcal{D}}(u^{0},\cdot)$ have slope less or equal 1. Therefore, using $|{-}\partial{\mathcal{D}}(u^{o},\cdot)|\leq 1$ the second estimate in (2.12) follows.
We can combine Propositions 2.8 and 2.9 and obtain the De Giorgi’s lemma for $\mathrm{gMGS}$ $(M,{\mathcal{E}},{\mathcal{D}},\psi)$, cf. also [RMS08, Lemma 4.5].
Theorem 2.10 (De Giorgi’s lemma for $\mathrm{gMGS}$)
Under Hypothesis 2.5, any measurable selection $(0,\sigma_{*})\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ fulfills
${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma\,\psi\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}+\int_{0}^{\sigma}\psi^{*}\big{(}\big{|}\partial{\mathcal{E}}\big{|}(\widetilde{u}_{\rho})\big{)}\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{o}).$ (2.13)
With the aim of improving (2.13) to an equality, taking into account the slope estimate, it is thus natural to check for cases in which $|\partial{\mathcal{D}}(u^{o}\cdot)|=1$. This is true for a geodesic distance.
@@
Suppose in addition that the metric space $(M,{\mathcal{D}})$ is a geodesic space. Then, for every $u^{o}\in M$ we have $|\partial{\mathcal{D}}(u^{o},\cdot)|(u)=1$ for every $u\neq u^{o}$.
Proof. Clearly, it suffices to show that $|\partial{\mathcal{D}}(u^{o},\cdot)|(u)\geq 1$. For this, let us consider the constant-speed geodesic $\gamma$ connecting $u^{0}$ to $u$, such that ${\mathcal{D}}(\gamma_{t},\gamma_{s})=(t{-}s){\mathcal{D}}(u^{o},u)$ for all $0\leq s\leq t\leq 1$. Then,
$|\partial{\mathcal{D}}(u^{o},\cdot)|(u)$ $=\limsup_{v\to u}\frac{({\mathcal{D}}(u^{o},u){-}{\mathcal{D}}(u^{o},v))^{+}}{{\mathcal{D}}(u,v)}$
$\geq\limsup_{t\to 0^{+}}\frac{({\mathcal{D}}(u^{o},u){-}{\mathcal{D}}(u^{o},\gamma_{t}))^{+}}{{\mathcal{D}}(u,\gamma_{t})}=\lim_{t\to 0^{+}}\frac{{\mathcal{D}}(u^{o},u)-t{\mathcal{D}}(u^{o},u)}{(1{-}t){\mathcal{D}}(u^{o},u)}=1\,.$
@@
(with $R>0$ a given constant), we have |(±$|\partial({\pm}{\mathcal{D}})(u^{o},\cdot)|(u)=0$ whenever $|u{-}u^{o}|>R$.
We next provide two examples showing that, without a geodesic metric and without a continuous slope, we cannot expect (2.13) to holds as an equality.
Example 2.12 ($(M,{\mathcal{D}})$ not geodesic)
@@
Starting with $u^{o}>1$ and setting $\sigma_{*}=\big{(}(u^{o})^{2}{-}1\big{)}^{-1/2}$ we obtain
$J_{\sigma}(u^{o})=\mathrm{argmin}\big{\{}\,\frac{1}{\sigma}{\mathcal{D}}(u^{o},u)^{2}+{\mathcal{E}}(u)\>\big{|}\>u\in{\mathbb{R}}\,\big{\}}=\begin{cases}\frac{u^{o}}{1+\sigma}&\text{for }\sigma<\sigma_{*},\\ \big{\{}\frac{u^{o}}{1+\sigma},0\big{\}}&\text{for }\sigma=\sigma_{*},\\ 0&\text{for }\sigma>\sigma_{*}.\end{cases}$
@@
Note that $|\partial{\mathcal{E}}|(\widetilde{u}_{\sigma})=|\partial{\mathcal{E}}|(0)=0$ for $\sigma>\sigma_{*}$ but $\psi^{\prime}\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}=\frac{1}{\sigma}{\mathcal{D}}(u^{o},0)=1/\sigma\gneqq 0$. Hence, the first estimate in Proposition 2.9 holds as an equality, but the second estimate is strict because of $\big{|}\partial({-}{\mathcal{D}})(u^{o},\cdot)\big{|}(0)=0<1$.
@@
Example 2.13 (Slope of ${\mathcal{E}}$ is not continuous)
@@
Starting at $u^{o}=1$ we find the unique variational interpolant $\widetilde{u}_{\sigma}=\max\{1{-}\sigma,0\}$. The curve $\sigma\mapsto\widetilde{u}_{\sigma}$ is absolutely continuous but the slope along the curve is discontinuous, namely
@@
This time the first estimate in Proposition 2.9 is strict for $\sigma\geq 1$, and hence (2.13) is also a strict inequality for $\sigma>1$.
2.2 Equality in De Giorgi’s estimate
The discussion in Section 2.1 has highlighted the link between two properties (one related to the geometry of the space, the other to the driving energy) and equality in the De Giorgi’s Lemma. With Theorem 2.14 we now prove that the joint validity of such properties is a sufficient, albeit rather strong, condition for a $\mathrm{gMGS}$ to guarantee that all measurable variational interpolants satisfy estimate (2.13) with equality.
@@ 2.2 Equality in De Giorgi’s estimate
    $(M,{\mathcal{D}})$ is a geodesic space
@@
    the mapping $u\mapsto|\partial{\mathcal{E}}|(u)$ is continuous and that there exists a continuous function $\omega:M{\times}M\to[0,1]$ with $\omega(u,u)=0$ for all $u\in M$ such that the following uniform slope estimate holds:
    $\forall\,u,w\in M:\quad{\mathcal{E}}(w)\geq{\mathcal{E}}(u)-|\partial{\mathcal{E}}|(u)\,{\mathcal{D}}(u,w)-\omega(u,w){\mathcal{D}}(u,w).$ (2.15)
@@
Then, for any measurable selection $(0,\sigma_{*})\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ the relation (2.13) holds with equality.
@@
Proof. It suffices to show that the slope estimate from Proposition 2.9 holds with equality and combine this with Lemma 2.11. Since the upper estimate $|\partial{\mathcal{E}}|(\widetilde{u}_{\sigma})\leq\psi^{\prime}\big{(}\frac{1}{\tau}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}$ is clear, it remains to show $|\partial{\mathcal{E}}|(\widetilde{u}_{\sigma})\geq\psi^{\prime}\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}$.
@@
We consider the geodesic curve $[0,1]\ni\theta\mapsto\gamma_{\theta}$ with $\gamma_{0}=u^{o}$ and $\gamma_{1}=\widetilde{u}_{\sigma}$. Since $\gamma_{1}=\widetilde{u}_{\sigma}$ is a global minimizer for $\Phi_{\sigma}(u^{o};\cdot)$ we can compare with $u=\gamma_{\theta}$ and obtain
${\mathcal{E}}(\gamma_{\theta})-{\mathcal{E}}(\widetilde{u}_{\sigma})$ $\geq\sigma\,\psi\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}-\sigma\,\psi\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\gamma_{\theta})\big{)}$
$\geq\psi^{\prime}\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\gamma_{\theta})\big{)}\,\big{(}{\mathcal{D}}(\gamma_{0},\gamma_{1})-{\mathcal{D}}(\gamma_{0},\gamma_{\theta})\big{)}=\psi^{\prime}\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\gamma_{\theta})\big{)}\,{\mathcal{D}}(\gamma_{\theta},\widetilde{u}_{\sigma}),$
@@
where we used the convexity of $\psi$ and the fact that $(\gamma_{\theta})_{\theta}$ is a geodesic.
Exploiting the slope condition (2.15) with $u=\gamma_{\theta}$ and $w=\widetilde{u}_{\sigma}$ we obtain
@@
$|\partial{\mathcal{E}}|(\gamma_{\theta})\geq\frac{{\mathcal{E}}(\gamma_{\theta})-{\mathcal{E}}(\widetilde{u}_{\sigma})}{{\mathcal{D}}(\gamma_{\theta},\widetilde{u}_{\sigma})}-\omega(\gamma_{\theta},\widetilde{u}_{\sigma})\geq\psi^{\prime}\big{(}\frac{1}{\tau}{\mathcal{D}}(u^{o},\gamma_{\theta})\big{)}-\omega(\gamma_{\theta},\widetilde{u}_{\sigma}).$
@@
We can now take the limit $\theta\nearrow 1$ and the continuity properties of the slope and $\psi^{\prime}$ yield the desired estimate $|\partial{\mathcal{E}}|(\widetilde{u}_{\sigma})\geq\psi^{\prime}\big{(}\frac{1}{\sigma}{\mathcal{D}}(u^{o},\widetilde{u}_{\sigma})\big{)}$.

### v3

@@ 2.1 Metric energy identity and the De Giorgi estimate: statements and examples
Under Hypothesis 2.5, any measurable selection $(0,\sigma_{*})\ni\sigma\to\widetilde{u}_{\sigma}\in J_{\sigma}(u^{\circ})$ fulfills
${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma\,\psi\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big)+\int_{0}^{\sigma}\psi^{*}\big(\psi^{\prime}\big(\frac{1}{\rho}\mathsf{d}^{\pm}_{\rho}(u^{\circ})\big)\big)\;\!\mathrm{d}\rho={\mathcal{E}}(u^{\circ}).$ (2.12)
There is a straightforward way to relate the metric energy identity to De Giorgi estimate, and that is throughout the following result in which the slope at $\widetilde{u}_{\sigma}$ is estimated in terms of the slope of the distance function ${\mathcal{D}}(u^{\circ},\cdot)$. In fact, (2.13) below extends to the setup of a $\mathrm{gMGS}$ $(M,{\mathcal{E}},{\mathcal{D}},\psi)$, the slope estimate proved in [AGS05, Lem. 3.1.3] in the quadratic case $\psi(r)=\frac{1}{2}r^{2}$. For completeness, we also record that a version of (2.13) was proved in [RMS08, Lemma 4.4] for non-differentiable dissipation potentials $\psi$. Note that this estimate is the metric counterpart of the Euler-Lagrange equation in the Banach-space setting, and hence it is less precise for a general $\mathrm{gMGS}$.
@@
$\big|\partial{\mathcal{E}}\big|(\widetilde{u}_{\sigma})\leq\psi^{\prime}\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big)\;\big|\partial({-}{\mathcal{D}})(u^{\circ},\cdot)\big|(\widetilde{u}_{\sigma})\leq\psi^{\prime}\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big).$ (2.13)
@@
${\mathcal{E}}(\widetilde{u}_{\sigma})-{\mathcal{E}}(v)$ $=\Phi_{\sigma}(\widetilde{u}_{\sigma})-\Phi_{\sigma}(v)-\sigma\psi\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big)+\psi\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},v)\big)$
$\stackrel{{\scriptstyle(1)}}{{\leq}}\sigma\Big(\psi\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},v)\big)-\psi\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big)\Big)$
$\stackrel{{\scriptstyle(2)}}{{\leq}}\psi^{\prime}\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big)\>\Big({-}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})+{\mathcal{D}}(u^{\circ},v)\Big).$
where (1) is due to $\widetilde{u}_{\sigma}\in J_{\sigma}$, whereas (2) follows by the convexity of $\psi$. Taking the positive part on both sides (using $\psi^{\prime}\geq 0$), dividing by ${\mathcal{D}}(\widetilde{u}_{\sigma},v)$, and taking the limsup for $v\to\widetilde{u}_{\sigma}$ gives the first estimate in (2.13).
By the triangle inequality for the distance $\mathcal{D}$}, for any fixed $u^{\circ}\in M$ the functions ${\mathcal{D}}(u^{\circ},\cdot)$ and ${\mathcal{D}}(u^{\circ},\cdot)$ have slope less or equal 1. Therefore, using $|{-}\partial{\mathcal{D}}(u^{\circ},\cdot)|\leq 1$, the second estimate in (2.13) follows.
We can combine Propositions 2.8 and 2.9 and obtain the De Giorgi lemma for $\mathrm{gMGS}$ $(M,{\mathcal{E}},{\mathcal{D}},\psi)$, cf. also [RMS08, Lemma 4.5].
Theorem 2.10 (De Giorgi lemma for $\mathrm{gMGS}$)
Under Hypothesis 2.5, any measurable selection $(0,\sigma_{*})\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}$ fulfills
${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma\,\psi\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big)+\int_{0}^{\sigma}\psi^{*}\big(\big|\partial{\mathcal{E}}\big|(\widetilde{u}_{\rho})\big)\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{\circ}).$ (2.14)
With the aim of improving (2.14) to an equality, taking into account the slope estimate, it is thus natural to check for cases in which $|\partial{\mathcal{D}}(u^{\circ},\cdot)|=1$. This is true for a geodesic distance.
@@
Suppose in addition that the metric space $(M,{\mathcal{D}})$ is a geodesic space. Then, for every $u^{\circ}\in M$ we have $|\partial{\mathcal{D}}(u^{\circ},\cdot)|(u)=1$ for every $u\neq u^{\circ}$.
Proof. Clearly, it suffices to show that $|\partial{\mathcal{D}}(u^{\circ},\cdot)|(u)\geq 1$. For this, let us consider the constant-speed geodesic $\gamma$ connecting $u^{\circ}$ to $u$, such that ${\mathcal{D}}(\gamma_{t},\gamma_{s})=(t{-}s){\mathcal{D}}(u^{\circ},u)$ for all $0\leq s\leq t\leq 1$. Then,
$|\partial{\mathcal{D}}(u^{\circ},\cdot)|(u)$ $=\limsup_{v\to u}\frac{({\mathcal{D}}(u^{\circ},u){-}{\mathcal{D}}(u^{\circ},v))^{+}}{{\mathcal{D}}(u,v)}$ (2.15)
$\geq\limsup_{t\to 0^{+}}\frac{({\mathcal{D}}(u^{\circ},u){-}{\mathcal{D}}(u^{\circ},\gamma_{t}))^{+}}{{\mathcal{D}}(u,\gamma_{t})}=\lim_{t\to 0^{+}}\frac{{\mathcal{D}}(u^{\circ},u)-t{\mathcal{D}}(u^{\circ},u)}{(1{-}t){\mathcal{D}}(u^{\circ},u)}=1\,.$
Remark 2.12
We highlight that, for the validity of the above result it is sufficient for $(M,{\mathcal{D}})$ to be just a length space [BBI01, Chap. 2], namely with the property that the distance ${\mathcal{D}}(u^{\circ},u)$ between any two points $u^{\circ},\,u\in M$ can be realized as the infimum of the lengths of the curves joining them. Then, for the lower estimate in (2.15) the role of the constant-speed geodesic $\gamma$ connecting $u^{\circ}$ to $u$ can be played by a curve with length arbitrarily close to ${\mathcal{D}}(u^{\circ},u)$.
@@
with $R>0$ a given constant (indeed, note that $({\mathbb{R}}^{n},{\mathcal{D}})$ is not a length space, either), we have $|\partial({\pm}{\mathcal{D}})(u^{\circ},\cdot)|(u)=0$ whenever $|u{-}u^{\circ}|>R$.
We next provide two examples showing that, without a geodesic metric and without a continuous slope, we cannot expect (2.14) to hold as an equality.
Example 2.13 (Failure of equality in (2.14) if $(M,{\mathcal{D}})$ not geodesic)
@@
Starting with $u^{\circ}>1$ and setting $\sigma_{*}=\big((u^{\circ})^{2}{-}1\big)^{-1/2}$ we obtain
$J_{\sigma}=\mathrm{Argmin}\big\{\,\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},u)^{2}+{\mathcal{E}}(u)\>\big|\>u\in{\mathbb{R}}\,\big\}=\begin{cases}\frac{u^{\circ}}{1+\sigma}&\text{for }\sigma<\sigma_{*},\\ \big\{\frac{u^{\circ}}{1+\sigma},0\big\}&\text{for }\sigma=\sigma_{*},\\ 0&\text{for }\sigma>\sigma_{*}.\end{cases}$
@@
Note that $|\partial{\mathcal{E}}|(\widetilde{u}_{\sigma})=|\partial{\mathcal{E}}|(0)=0$ for $\sigma>\sigma_{*}$ but $\psi^{\prime}\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big)=\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},0)=1/\sigma\gneqq 0$. Hence, the first estimate in Proposition 2.9 holds as an equality, but the second estimate is strict because of $\big|\partial({-}{\mathcal{D}})(u^{\circ},\cdot)\big|(0)=0<1$.
@@
Example 2.14 (Failure of equality in (2.14) if the slope of $\mathcal{E}$} is not continuous)
@@
Starting at $u^{\circ}=1$ we find the unique variational interpolant $\widetilde{u}_{\sigma}=\max\{1{-}\sigma,0\}$. The curve $\sigma\mapsto\widetilde{u}_{\sigma}$ is absolutely continuous but the slope along the curve is discontinuous, namely
@@
This time, the first estimate in Proposition 2.9 is strict for $\sigma\geq 1$, and hence (2.14) is also a strict inequality for $\sigma>1$.
2.2 Equality in the De Giorgi estimate
The discussion in Section 2.1 has highlighted the link between two properties (one related to the geometry of the space, the other to the driving energy and its slope), and equality in the De Giorgi Lemma. With Theorem 2.15 we now prove that the joint validity of such properties is a sufficient, albeit rather strong, condition for a $\mathrm{gMGS}$ to guarantee that all measurable variational interpolants satisfy estimate (2.14) with equality. The idea is that, if $|\partial{\mathcal{E}}|$ is upper semicontinuous, points where the slope is ‘too big’ are avoided by minimizers.
@@ 2.2 Equality in the De Giorgi estimate
    $(M,{\mathcal{D}})$ is a geodesic space, and
@@
    $\mathcal{E}$} is continuous on $M$,
  3. 3.
    $|\partial{\mathcal{E}}|$ is upper semicontinuous on $M$.
@@
Then, for any measurable selection $(0,\sigma_{*})\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}$ the relation (2.14) holds with equality.
@@
Proof. It suffices to show that the slope estimate from Proposition 2.9 holds with equality and combine this with Lemma 2.11. Since the upper estimate $|\partial{\mathcal{E}}|(\widetilde{u}_{\sigma})\leq\psi^{\prime}\big(\frac{1}{\tau}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big)$ is clear, it remains to show $|\partial{\mathcal{E}}|(\widetilde{u}_{\sigma})\geq\psi^{\prime}\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big)$.
@@
We consider the geodesic curve $0,1]\ni\theta\mapsto\gamma_{\theta}$ with $\gamma_{0}=u^{\circ}$ and $\gamma_{1}=\widetilde{u}_{\sigma}$. Since $\gamma_{1}=\widetilde{u}_{\sigma}$ is a global minimizer for $\Phi_{\sigma}(\cdot)$, we can compare with $u=\gamma_{\theta}$ and obtain
${\mathcal{E}}(\gamma_{\theta})-{\mathcal{E}}(\gamma_{1})$ $={\mathcal{E}}(\gamma_{\theta})-{\mathcal{E}}(\widetilde{u}_{\sigma})\geq\sigma\,\psi\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big)-\sigma\,\psi\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\gamma_{\theta})\big)$
$\geq\psi^{\prime}\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\gamma_{\theta})\big)\,\big({\mathcal{D}}(\gamma_{0},\gamma_{1})-{\mathcal{D}}(\gamma_{0},\gamma_{\theta})\big)=\psi^{\prime}\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\gamma_{\theta})\big)\,{\mathcal{D}}(\gamma_{\theta},\gamma_{1}),$
@@
where we used the convexity of $\psi$ and the fact that $(\gamma_{\theta})_{\theta}$ is a geodesic. Thus, we have
@@
$S_{\theta}:=\frac{{\mathcal{E}}(\gamma_{\theta}){-}{\mathcal{E}}(\gamma_{1})}{{\mathcal{D}}(\gamma_{\theta},\gamma_{1})}\geq\psi^{\prime}\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\gamma_{\theta})\big)\text{ \ for all }\theta\in[0,1).$
Now fix $\theta$ and define the function $h_{\theta}(s)={\mathcal{E}}(\gamma_{s})-S_{\theta}{\mathcal{D}}(\gamma_{s},\gamma_{1})$ for $s\in[\theta,1]$. Since $\mathcal{E}$} is continuous, $h_{\theta}$ is continuous, too, and satisfies $h_{\theta}(\theta)=h_{\theta}(1)={\mathcal{E}}(\gamma_{1})$. Hence, there exists a maximizer $t_{\theta}\in(\theta,1)$ of $h_{\theta}$. For this $t_{\theta}$ and $s\in(t_{\theta},1]$ we have, by construction,
${\mathcal{E}}(\gamma_{t_{\theta}})-{\mathcal{E}}(\gamma_{s})\geq S_{\theta}({\mathcal{D}}(\gamma_{t_{\theta}},\gamma_{1})){-}{\mathcal{D}}(\gamma_{s},\gamma_{1})\big)=S_{\theta}{\mathcal{D}}(\gamma_{t_{\theta}},\gamma_{s}).$
Dividing by ${\mathcal{D}}(\gamma_{t_{\theta}},\gamma_{s})$ and taking the limit $s\searrow t_{\theta}$ we find
$|\partial{\mathcal{E}}|(\gamma_{t_{\theta}})\geq S_{\theta}\geq\psi^{\prime}\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\gamma_{\theta})\big).$
In the limit $\theta\nearrow 1$ we have $\gamma_{t_{\theta}}\to\gamma_{1}=\widetilde{u}_{\sigma}$, and the upper semicontinuity of $|\partial{\mathcal{E}}|$ yields
@@
$|\partial{\mathcal{E}}|(\widetilde{u}_{\sigma})=|\partial{\mathcal{E}}|(\gamma_{1})\geq\limsup_{\theta\nearrow 1}|\partial{\mathcal{E}}|(\gamma_{t_{\theta}})\geq\lim_{\theta\nearrow 1}\psi^{\prime}\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\gamma_{\theta})\big)=\psi^{\prime}\big(\frac{1}{\sigma}{\mathcal{D}}(u^{\circ},\widetilde{u}_{\sigma})\big),$
which is the desired opposite inequality to (2.12), and the assertion follows.